Two people are pushing a stalled car; the mass of the car is 1850 kg. one person applies a force of 275 N to the car, while the other applies a force of 395 N. Both forces act in the same direction. A third force of 560N also acts on the car, but in a direction opposite to that in which the people are pushing. This force arises because of friction and the extent to which the pavement opposes the motion of the tires. Find the acceleration of the car.


1
Expert's answer
2021-01-26T08:17:12-0500

As we know from Newton's second law, the acceleration is net force divided by mass. The net force is a vector sum of all forces, so:


"F_\text{net}=275+395-560=110\text{ N},\\\space\\
a=\frac{F_\text{net}}{m}=\frac{110}{1850}=0.059\text{ m/s}^2."